Solve for h
Tan 75=(h/510-h)
Assignment due by 7:15 today,need immediate help if possible. Thank you

Answer by Gogonati(809)

Solution: Assume that 510-h different from zero, multiply both sides by 510-h,
75(510-h)=h => 38250-75h=h =>38250=76h => h=38250/76 => h=19125/38.
Done.